OpenStudy (anonymous):

Write an equation in slope-point form for the line that passes through A(1, 4) and B(6, 8). Then write in slope-intercept form. Can someone help me with this?

OpenStudy (zale101):

@AngelWu have you tried solving the problem? :)

OpenStudy (anonymous):

@Zale101 I first tried to find the slope from the two points given and I got 4/5 but I don't know how to write it into slope-point form

OpenStudy (texaschic101):

slope = (8 - 4) / (6 - 1) slope(m) = 4/5 well....you got the slope right. Now use point slope formula: y - y1 = m(x - x1) slope(m) = 4/5 use either of your points...(1,4)...x1 = 1 and y1 = 4 now sub y - 1 = 4/5(x - 4) <=== point slope form Now we need slope intercept form y = mx + b y - 1 = 4/5(x - 4) --- distribute through the parenthesis y - 1 = 4/5x - 16/5 -- add 1 to both sides y = 4/5x - 16/5 + 1 y = 4/5x - 16/5 + 5/5 y = 4/5x - 11/5 <=== slope intercept form

OpenStudy (anonymous):

Thank you! @texaschic101